OpenStudy (anonymous):

Identify the volume and surface area of the sphere in terms of π. Note; The sphere has a radius of 9cm.

OpenStudy (anonymous):

the answer is 972.49

OpenStudy (anonymous):

v=972.49

OpenStudy (anonymous):

surface area is 324.17

OpenStudy (anonymous):

do you understand?

OpenStudy (anonymous):

Thank you! I somehow came up with 228,906.

OpenStudy (anonymous):

and there is also pi after the answer

OpenStudy (anonymous):

is that all?

OpenStudy (anonymous):

yeah because I didn't know what to do with the 3/4 you divide it and get something like 0.666665, but I don't know what to do with that with the formula afterwards.

OpenStudy (anonymous):

so do you need the formula

Directrix (directrix):

Answers have to be in terms of >>in terms of π.

OpenStudy (anonymous):

Yeah, it's alright. I won't be good at maths like I used to be. Ever. It's just what happens when you end up with seizures and no medication. Kinda hard on the brain to be able to do any problems.

Directrix (directrix):

I did not get that volume. Let me check my work.

Directrix (directrix):

@joojoe did not show work so I'll start from scratch.

OpenStudy (anonymous):

sorry about that, ok wait

OpenStudy (anonymous):

A=4πr2

OpenStudy (anonymous):

V=4/3πr3

OpenStudy (anonymous):

r=9 cm \[Volume ~of~sphere~V=\frac{ 4 }{ 3 } \pi \times r^3=\frac{ 4 }{ 3 } \pi \times 9^3=?\] \[Surface~area ~A=4 \pi r^2=4 \pi \times 9^2=?\]

OpenStudy (anonymous):

bye if you have any question let me know i would be happy to help

OpenStudy (anonymous):

I got 101,736 for area with that formula.

OpenStudy (anonymous):

I'm probably wrong..

OpenStudy (anonymous):

Options are: V = 243π cm3; S = 162π cm2 V = 972π cm3; S = 162π cm2 V = 243π cm3; S = 324π cm2 V = 972π cm3; S = 324π cm2

OpenStudy (anonymous):

\[A=4~\pi \times 81=324 \pi~ cm^2\]

OpenStudy (anonymous):

calculate V

OpenStudy (anonymous):

3,053

OpenStudy (anonymous):

\[V=4 \pi \times 3 \times 9 \times 9=? cm^3\]

OpenStudy (anonymous):

Okay. 33,912 so far, holdon

OpenStudy (anonymous):

that working is wrong

OpenStudy (anonymous):

305,208 is what I got for the final answer with the formula.

OpenStudy (anonymous):

\[V=972 \pi ~cm^3\]

OpenStudy (anonymous):

\[v=4/3\Pi r ^{3}\]

OpenStudy (anonymous):

do not substitute the value of pi,write it as pi.

OpenStudy (anonymous):

yes she know that.

OpenStudy (anonymous):

I'm just awful at this, and getting really seizureish, don't like having to try to do it all on my notebook paper.

OpenStudy (anonymous):

Just to get it all wrong.

OpenStudy (anonymous):

@Calamitylikespie should i write it for you

OpenStudy (anonymous):

Getting me teary eyed, and frustrated, I'm to vulnerable to this considering how awful I do at it.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

It'd maybe help out.

OpenStudy (anonymous):